a total of 144 votes were cast for student body president . If there were twice as many votes for cindy as there were for Jim, how many votes did Cindy receive ?
Answer by ikleyn(52944)

.

Cindy + Jim = 144


Cindy = 2*Jim


2*Jim + Jim = 144


3*Jim = 144


  Jim = 144/3 = 48.


Cindy = 2*Jim = 2*48 = 96.    ANSWER

Solved.